A room is 20 feet long. If a couch that is 12 1/3 feet long is to be centered in the room, how big of a table can be placed on either side of the couch?
I get that you would have to subtract the 12 1-3 from the 20 feet. I then get you have to take that answer and divide it by 2 (because the couch is centered so you would have the amount of space on either side). The problem I am having is getting the fraction to come out right and showing my work correctly.

Answer by rvfer(117)

20-12 1/3=7 2/3
(7 2/3)/2=
(23/3)/(2/1)=
23/3*1/2=23/6
reduce 23/6=3 5/6ft each side